OpenStudy (anonymous):

Finding product. (-5a^2)^3·a^5 . i tried and i got 8.3 a^6, and i got it wrong:( HELP PLEASEEEE!!

OpenStudy (lgbasallote):

\[(-5a^2)^3 \implies -5^3 \times (a^2)^3\] start tehre

OpenStudy (anonymous):

ya so I got -5^3a^5??

OpenStudy (lgbasallote):

firstly.. \[(a^2)^3 \implies a^{2\times 3}\] second..what's -5^3? simplify it

OpenStudy (anonymous):

-125

OpenStudy (anonymous):

-125a^9??

OpenStudy (anonymous):

is that a ^3x3??

OpenStudy (lgbasallote):

what is 2 x 3?

OpenStudy (lgbasallote):

\[\huge (a^2)^3 \implies a^{2\times 3}\]

OpenStudy (anonymous):

oh i thought it was 3x3.. so is it -125a^6?

OpenStudy (lgbasallote):

yup

OpenStudy (lgbasallote):

now multiply to a^5 \[\huge -125a^6 \times a^5 \implies -125a^{6+5}\]

OpenStudy (anonymous):

-125a^11

OpenStudy (anonymous):

THANK YOU!!

OpenStudy (lgbasallote):

you're welcome ^_^